# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

### v1

@@ 3.1 POPCON for tandem mirrors
To generate tandem mirror POPCONs, the steady-state 0-D tandem mirror transport equations based on those found in Mirin et al. (1983); Hua & Fowler (2004) are taken to the steady state limit and all derivatives are taken to zero. However, consistent with the typical POPCON formulation, radial variation in plasma parameters (including radial variation in axial confinement) has been allowed. It is assumed that the tandem mirror is operating a classical tandem without thermal barriers (Fowler & Logan, 1977). Under this assumption electrons in the central cell and end plug are in thermal contact and $T_{e}\equiv T_{ec}=T_{ep}$. This conservative assumption dramatically simplifies the system of equations which must be solved to make a POPCON. Under these assumptions, the primary energy and particle balance equations in the POPCON system are:
@@
where subscripts $c$ and $p$ denote central cell and end plug quantities respectively, subscripts $i$ and $e$ denote electrons and ions respectively, $n_{c}\equiv n_{ec}=n_{ic}$ is the density assuming quasineutrality, $T_{s}$ is the temperature of species $s$, $\ell_{c}$ is the central cell length, $a_{c}$ is the central cell radius, $\Gamma_{\text{gas}}$ is the gas fueling rate, $\tau_{c}\approx\tau_{p}\approx\tau_{E}$ is the confinement time, $\mathcal{P}_{ie}$ is the collisional power density transfer from ions to electrons, $P_{\mathrm{RF},s}$ is the applied RF heating power to the central cell, $\mathcal{P}_{\mathrm{fus},s}$ is the fusion power density from alpha slowing down applied to species $s$, and $\mathcal{P}_{\mathrm{rad}}$ is the power density radiated by bremsstrahlung and synchrotron radiation.
In order to create a set of POPCONs, (13) is solved for heating $P_{\mathrm{RF}}$ (with some fraction directed at ions and electrons respectively) and fueling $\Gamma_{\mathrm{gas}}$ at a given operating point with fixed $T_{ic}$ and $n_{c}$. To solve this system, all variables must be written as a function of $T_{ic}$ and $n_{c}$ and fixed end plug and central cell parameters: $B_{m}$, $a_{m}$, $n_{p}$, $\ell_{c}$, $B_{0c}$ ($T_{e}$ may be solved for as a dependent variable). This is done by applying the following definitions; the confinement time due to classical radial and axial transport (the dominant transport modes expected in the tandem mirror) $\tau_{c}$ is defined:
@@
where the Pastukhov confinement of ions in the central cell by the end plug potential is defined (Pastukhov, 1974; Cohen et al., 1978):
@@
$\tau_{\rho}=\left(\frac{a_{c}}{\rho_{ic}}\right)^{2}\tau_{ii},$ (19)
where $\rho_{ic}=v_{\mathrm{th}ic}/\Omega_{ic}$ is the central cell ion Larmor radius. Returning to (13), the electron confining potential in the plug can be defined for classical tandem operation with electron-drag dominated end plugs by solving transcendental equation (Fowler & Logan, 1977; Hua & Fowler, 2004):
@@
Using this POPCON power-balance, it was possible to reproduce the tandem mirror operating points identified in Fowler et al. (2017) when the same assumptions were made i.e. no classical transport (19), fusion heating (22), or radiation (24), and all central cell RF heating was taken to be electron directed. The $T_{ic}=150~{}\mathrm{keV}$, $\beta_{c}\sim 1$ operating points described in Fowler et al. (2017) were found to within a small factor of error (<1.25) in $T_{i}$ that was determined to be associated with a discrepancy in the fusion cross sections used in their model and the Bosch-Hale cross sections used here, they considered $T_{ic}\equiv T_{ec}$ when in reality $T_{ic}\sim T_{ec}$ at these parameters, as well as the sometimes poor convergence of the model here at very high values of $\beta_{c}$. When the additional terms used here were included it was found that the operating points in Fowler et al. (2017) could not be attained. Radiation and classical transport significantly reduced performance (both terms become large at the very high $\beta\sim 1$ operating points considered there, though, it should be noted that the expression for classical transport used here is not valid when $\beta\sim 1$).
@@ 3.2 Identifying operating points with model end plugs
$a_{m}$ [m] $B_{m}$ [T] $B_{0c}$ [T] $\ell_{c}$ [m] $\langle n_{p}\rangle$ [m3]
0.10 25.0 5.0 50.0 $2.0\times 10^{20}$
0.15 25.0 3.0 50.0 $1.5\times 10^{20}$
0.20 25.0 2.6 50.0 $1.25\times 10^{20}$
0.25 25.0 2.4 50.0 $1.0\times 10^{20}$
Table 1: Tandem mirror parameters used in the POPCONs which appear in Figure04
Figure 4: POPCON plots for tandem mirrors at four different plasma radii through the mirror throat, (a) 0.1 m, (b) 0.15 m, (c) 0.2 m, and (d) 0.25 m. The other parameters used are shown in Table 1. Operating points are shown as a function of central cell density $n_{c}$ divided by end plug density $n_{p}$ versus central cell ion temperature $T_{i}$. Blue and red filled contours are of heating power to the central cell required at a given ($n_{c}$, $T_{i}$) operating point (blue regions indicate ignition). Regions which are off the scale for positive or negative and are inaccessible are denoted with red and blue hatching respectively. Black contour lines are of fusion power in the central cell. Green contour lines are of central cell vacuum $\beta$.
With the components of the POPCON system of equations (13) defined, it is possible to solve for the required heating and fueling at each location in ($n_{c}$, $T_{ic}$) space for a given set of end plug radial density profiles and engineering parameters. POPCONs with simplified end plugs were solved to establish target parameters for the end plug to be simulated with RealTwin. In these simplified end plugs, the density profile was taken to be flat, the ion directed fraction of $P_{\mathrm{RF}}$ was taken to be $80\%$, and the end plug and central cell parameters shown in Table 1 were used. End plug parameters were taken to be fixed besides the plasma radius through the mirror throat $a_{m}$ and the average density within the plug $\langle n\rangle$ which were varied to determine a target operating range. The central field was increased until either the number of alpha gyroradii across the plasma radius $N_{\alpha}>5$ defined using:
@@
was obtained or an ignited operating point with $P_{\mathrm{fus}}>150~{}\mathrm{MW}$ and $\beta<0.6$ was obtained. Average end plug density $\langle n_{p}\rangle$ was varied between $1.0\times 10^{20}~{}\mathrm{m}^{-3}$ and $2.0\times 10^{20}~{}\mathrm{m}^{-3}$.
The results of the POPCON analysis described here are shown in Figure 4 for the parameters defined in Table 1. POPCON analysis indicates that, for what will be shown later on are fairly reasonable end plug parameters, it is possible to create a classical tandem mirror with an ignited central cell. Large ignition regions are present in all cases when the plasma radius at the mirror throat is larger than 0.15 m. These results indicate that a tandem mirror pilot plant designed with these principles should easily be able to satisfy the NASEM fusion pilot plant requirements at lengths under 100 m. Using simplified equation for the net electric power:
@@
Assuming neutral beam efficiencies $\eta_{\mathrm{NBI}}=60\%$, an electrical conversion efficiency $\eta_{\mathrm{ele}}=50\%$ assuming the use of a Brayton cycle (Schleicher et al., 2001), a blanket neutron energy multiplication factor $C_{\mathrm{mult}}=1.1$, and $P_{\mathrm{NBI}}=30~{}\mathrm{MW}$ into the end plugs; it was found a tandem mirror would need to produce $157.4~{}\mathrm{MW}$ of power to satisfy the NASEM $50~{}\mathrm{MWe}$ requirement. Such an operating point is achievable with $B_{m}=25~{}\mathrm{T}$, $a_{m}=0.15~{}\mathrm{m}$, $\langle n\rangle_{p}=1.5\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=50~{}\mathrm{m}$, $\beta_{c}\cong 0.6$, $n_{c}/n_{p}\cong 0.55$, $T_{i}\cong 45~{}\mathrm{keV}$, and $T_{e}\cong 125~{}\mathrm{keV}$ based on this POPCON analysis. The fusion power density of this case is $\sim 3.5~{}\mathrm{MWm}^{-3}$ which is competitive with high-field tokamaks (Frank et al., 2022b). Performance of such a system could be further enhanced with technologies such as direct electrical conversion (Moir & Barr, 1973) or more efficient NBI utilizing advanced neutralization techniques and energy recovery (Fumelli et al., 1989; Grisham, 2009; Kovari & Crowley, 2010). It is noteworthy that this level of performance is achievable despite using generally more conservative parameters than the DT classical tandem proposed in Fowler et al. (2017) which had: $B_{m}=24~{}\mathrm{T}$, $a_{m}=0.21~{}\mathrm{m}$, $\langle n_{p}\rangle=2.6\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=55~{}\mathrm{m}$, $\beta_{c}\sim 1.0$, $n_{c}/n_{p}\sim 0.35$, $T_{i}=150~{}\mathrm{keV}$, and $T_{e}=150~{}\mathrm{keV}$.
@@
Figure 5: Contours of $T_{e}$ for operating points at given $\langle n\rangle$ and $E_{\mathrm{NBI}}$ satisfying the constraint equations (34). This plot uses the $a_{m}=0.15$ m POPCONs case found in Figure 4b.
@@
This POPCONs analysis exposed a number of surprising features of fusion relevant tandem mirror operation that were not previously well known. These discoveries are in part because, outside of the analysis in Fowler et al. (2017), axisymmetric tandem mirrors with high performance HTS magnets and high $R_{m}$ had never been considered. Older studies of tandem mirrors generally assumed end plug mirror ratios of order unity to account for minimum-B coil geometries and lower peak $B_{m}$ (Fowler & Logan, 1977). This analysis builds upon that in Fowler et al. (2017), and very importantly, it includes the effect of fusion alpha particle heating on the electron temperature. The key result here is that, when high-field mirror coils are available, classical tandem central cells can and should be operated at much higher density fractions $n_{c}/n_{p}$ than previously assumed. Much of the previous classical tandem mirror literature used $n_{c}/n_{p}=0.05-0.35$ (Dimov et al., 1976; Fowler & Logan, 1977; Hua & Fowler, 2004; Fowler et al., 2017). The analysis here suggests this value should be higher, $n_{c}/n_{p}=0.35-0.6$. Strong alpha heating of the electrons obtained by, increasing the central cell density, boosting confinement with large $R_{m}$ from axisymmetrization, and directly heating the central cell ions with RF to enhance the fusion rate, allows the central cell to reach electron temperatures $\sim 2-3~{}T_{i}$ without supplemental electron heating as shown in Figure 5. This operational mode takes advantage of tandem mirror confinement (15) improving more rapidly with increasing temperature than it degrades with increasing the density ratio.
@@
In conclusion, this scoping shows that for high-field axisymmetric mirrors there classical tandems are a valid alternative to the traditional thermal-barriered tandem operation (Baldwin & Logan, 1979) which can still reach appreciable Q in agreement with similar analyses by Pratt & Horton (2006) and Fowler et al. (2017). However, it has been found here that higher density operation than these previous analyses is possible thanks to alpha heating, somewhat relaxing the end plug performance criteria, but there are still some important caveats. The tandem mirror POPCON approach used here, much like the tokamak POPCON approach, is only able to identify operating points. It does not guarantee that an operating point it identifies is stable or achievable in actual operation. To determine this, a time dependent simulation which takes the mirror from an initial breakdown plasma to its final operating point must be performed. It is possible that, to actually reach the operational conditions proposed here, a transient heating source, such as a high-power electron cyclotron heating "sparkplug", might be needed to boost $T_{e}$ initially. This would drive up the end plug potentials and improve confinement enough for the central cell to generate sufficient alpha heating power to reach ignition. Furthermore, the effects of spatial gradients or turbulent transport were not included in this analysis. Generally, the spatial density gradients in tandem mirrors are predicted to be small (Mirin et al., 1983; Hua & Fowler, 2004), and the effects of electron temperature gradient turbulence, the main form of turbulent energy transport predicted to be present in tandem mirrors, are anticipated to be more tolerable than the temperature gradient driven turbulent transport in tokamaks (Hua & Fowler, 2004; Pratt & Horton, 2006). These topics will be treated in future work where gyrokinetic turbulence simulations at mirror central cell conditions will be performed to evaluate the growth rates of gradient driven instabilities. A time dependent solver based on those found in Mirin et al. (1983) and Hua & Fowler (2004) in order to evaluate the accessibility of these operating points under more realistic assumptions will also be implemented. In addition to these confinement issues, macrostability of the tandem system, specifically its stability against trapped particle modes and MHD, have not been addressed here and will be the topic of a future publication.
@@
For tandem mirror parameters similar to the ignited operating points in Figure 4b: $\tau_{c}=5~{}\mathrm{s}$, $\phi_{e}=400~{}\mathrm{keV}$, $\ell_{c}=50~{}\mathrm{m}$, $R_{mp}=8$, $R_{mc}=13.3$ (accounting for finite $\beta\sim 0.6$ in both mirror ratios), $\theta_{\mathrm{NBI}}=45^{\circ}$, $E_{i}=500~{}\mathrm{keV}$, $\mu=2.5~{}\mathrm{AMU}$, and $a_{m}\sqrt{R_{mp}}/\rho_{i0}=25$, it is calculated that $n_{c}/n_{p}>1000$ indicating DCLC likely cannot be stabilized with central cell outflow in the tandem mirrors under consideration here. As $\Gamma_{c,\mathrm{out}}$ is insufficient for DCLC stabilization, DCLC stability in Section 4 will be ensured by the suppression of the diamagnetic drift wave and the use of sloshing ions.
@@
4 Simulation of an optimized end plug using RealTwin
The POPCON technique outlined in the previous section has defined what parameters the end plug must achieve in order to obtain a good central cell: the end plug mirror coils must have the highest field and bore possible, and an end plug must achieve the highest plasma density possible for the fewest watts of input power. In this section an end plug will be optimized, first some assumptions will be used to simplify the problem and physics and engineering constraints will be imposed on the end plug. Next, those constraints will be used to perform an initial optimization of the plug and identify an operating region. Finally, simulations with the RealTwin will be performed to provide a high-fidelity optimization based on the prior constraint based optimization.
4.1 Assumptions and constraints

### v2

@@ 3.1 POPCON for tandem mirrors
To generate tandem mirror POPCONs, the steady-state 0-D tandem mirror transport equations based on those found in Mirin et al. (1983); Hua & Fowler (2004) are taken to the steady state limit and all derivatives are taken to zero. However, consistent with the typical POPCON formulation, radial variation in plasma parameters (including radial variation in axial confinement) has been allowed. It is assumed that the tandem mirror is operating a classical tandem without thermal barriers (Fowler & Logan, 1977). Under this assumption electrons in the central cell and end plug are in thermal contact and $T_{e}\equiv T_{ec}=T_{ep}$. This assumption dramatically simplifies the system of equations which must be solved to make a POPCON. Under these assumptions, the primary energy and particle balance equations in the POPCON system are:
@@
where subscripts $c$ and $p$ denote central cell and end plug quantities respectively, subscripts $i$ and $e$ denote electrons and ions respectively, $n_{c}\equiv n_{ec}=n_{ic}$ is the density assuming quasineutrality, $T_{s}$ is the temperature of species $s$, $\ell_{c}$ is the central cell length, $a_{c}$ is the central cell radius, $\Gamma_{\text{gas}}$ is the gas fuelling rate, $\tau_{c}\approx\tau_{p}\approx\tau_{E}$ is the confinement time, $\mathcal{P}_{ie}$ is the collisional power density transfer from ions to electrons, $P_{\mathrm{RF},s}$ is the applied RF heating power to the central cell, $\mathcal{P}_{\mathrm{fus},s}$ is the fusion power density from alpha slowing down applied to species $s$, and $\mathcal{P}_{\mathrm{rad}}$ is the power density radiated by bremsstrahlung and synchrotron radiation.
In order to create a set of POPCONs, (15) is solved for heating $P_{\mathrm{RF}}$ (with some fraction directed at ions and electrons respectively) and fuelling $\Gamma_{\mathrm{gas}}$ at a given operating point with fixed $T_{ic}$ and $n_{c}$. To solve this system, all variables must be written as a function of $T_{ic}$ and $n_{c}$ and fixed end plug and central cell parameters: $B_{m}$, $a_{m}$, $n_{p}$, $\ell_{c}$, $B_{0c}$ ($T_{e}$ may be solved for as a dependent variable). This is done by applying the following definitions; the confinement time due to classical radial and axial transport (the dominant transport modes expected in the tandem mirror) $\tau_{c}$ is defined:
@@
where the ambipolar confinement of ions in the central cell by the end plug potential is defined (Pastukhov, 1974; Cohen et al., 1978):
@@
$\tau_{\rho}=0.25\left(\frac{a_{c}}{\rho_{ic}}\right)^{2}\tau_{ii},$ (21)
where $\rho_{ic}=v_{\mathrm{th}ic}/\Omega_{ic}$ is the central cell ion Larmor radius. Returning to (15), the electron confining potential in the plug can be defined for classical tandem operation by solving the transcendental equation (Fowler & Logan, 1977; Hua & Fowler, 2004):
@@
Using this POPCON power-balance, it was possible to reproduce the tandem mirror operating points identified in Fowler et al. (2017) when the same assumptions were made i.e. no classical transport (21), fusion heating (24), or radiation (26), and all central cell RF heating was taken to be electron directed. The $T_{ic}=150~{}\mathrm{keV}$, $\beta_{c}\sim 1$ operating points described in Fowler et al. (2017) were found to within a small factor of error in $T_{i}$ (<1.25) that was determined to be associated with a discrepancy in the fusion cross sections used in their model and the Bosch-Hale cross sections used here, they considered $T_{ic}\equiv T_{ec}$ when in reality $T_{ic}\sim T_{ec}$ at these parameters, as well as the sometimes poor convergence of root-finding algorithm used to evaluate the model here at very high values of $\beta_{c}$. When the additional terms used here were included it was found that the operating points in Fowler et al. (2017) could not be attained. Radiation and classical transport significantly reduced performance (both terms become large at the very high $\beta\sim 1$ operating points considered there, though, it should be noted that the expression for classical transport used here is not valid when $\beta\sim 1$ and Fowler et al. (2017) assumed very high wall reflectivities $R_{w}$).
@@ 3.2 Identifying operating points with model end plugs
Figure 4: POPCON plots for tandem mirrors with $\ell_{c}=50~{}\mathrm{m}$ at four different plasma radii through the mirror throat, (a) 0.1 m, (b) 0.15 m, (c) 0.2 m, and (d) 0.25 m using $n_{p}=1.5\times 10^{20}~{}\mathrm{m}^{-3}$ and $B_{0c}=3.125~{}\mathrm{T}$ in all cases except for (a) where $B_{0c}=5.0~{}\mathrm{T}$. Operating points are shown as a function of central cell density $n_{c}$ divided by end plug density $n_{p}$ versus central cell ion temperature $T_{i}$. Blue and red filled contours are of heating power to the central cell required at a given ($n_{c}$, $T_{i}$) operating point (blue regions indicate ignition). Regions which are off the scale for positive or negative and are inaccessible are denoted with red and blue hatching respectively. Black contour lines are fusion power in the central cell. Green contour lines are central cell vacuum $\beta$.
With the components of the POPCON system of equations (15) defined, it is possible to solve for the required heating and fuelling at each location in ($n_{c}$, $T_{ic}$) space for a given set of end plug radial density profiles and engineering parameters. POPCONs with simplified end plugs were solved to establish target parameters for the end plug to be simulated with RealTwin. In these simplified end plugs, the density profile was taken to be flat, the ion directed fraction of $P_{\mathrm{RF}}$ was taken to be $80\%$. End plug parameters were taken to be fixed besides the plasma radius through the mirror throat $a_{m}$. The central cell field was increased until either the number of alpha gyroradii across the plasma radius $N_{\alpha}>5$ defined using:
@@
was obtained or the central cell vacuum mirror ratio reached 8 (this limit was imposed in the interest of simplifying control of central cell interchange driven by trapped particle modes based on the results in (Gerver et al., 1989)). Average end plug density was varied until ignited operating points appeared. However, for illustrative purposes the POPCONs in this section all have the same density.
The results of the POPCON analysis described here are shown in Figure 4. POPCON analysis indicates that, for what will be shown later on are fairly reasonable end plug parameters, it is possible to create a classical tandem mirror with an ignited central cell. Ignition regions are present in all cases when the plasma radius at the mirror throat is larger than 0.15 m. These results indicate that a tandem mirror pilot plant designed with these principles should be able to satisfy the NASEM fusion pilot plant requirements at lengths under 100 m. Using simplified equation for the net electric power:
@@
Assuming neutral beam efficiencies $\eta_{\mathrm{NBI}}=60\%$, an electrical conversion efficiency $\eta_{\mathrm{ele}}=50\%$ assuming the use of a Brayton cycle (Schleicher et al., 2001), a blanket neutron energy multiplication factor $C_{\mathrm{mult}}=1.1$, and $P_{\mathrm{NBI}}=30~{}\mathrm{MW}$ into the end plugs; it was found a tandem mirror would need to produce $157.4~{}\mathrm{MW}$ of power to satisfy the NASEM $50~{}\mathrm{MWe}$ requirement. Such an operating point is achievable with $B_{m}=25~{}\mathrm{T}$, $a_{m}=0.15~{}\mathrm{m}$, $\langle n\rangle_{p}=1.5\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=50~{}\mathrm{m}$, $\beta_{c}\cong 0.6$, $n_{c}/n_{p}\cong 0.55$, $T_{i}\cong 45~{}\mathrm{keV}$, and $T_{e}\cong 125~{}\mathrm{keV}$ based on this POPCON analysis. The fusion power density of this case is $\sim 3.5~{}\mathrm{MWm}^{-3}$ which is competitive with high-field tokamaks (Frank et al., 2022b). Performance of such a system could be further enhanced by large $a_{m}$ or with technologies such as direct electrical conversion (Moir & Barr, 1973) or more efficient NBI utilizing advanced neutralization techniques and energy recovery (Fumelli et al., 1989; Grisham, 2009; Kovari & Crowley, 2010). Values of $a_{m}<15~{}\mathrm{cm}$ have difficulty obtaining good performance without very high plug densities $\langle n_{p}\rangle\gg 2.0\times 10^{20}~{}\textrm{m}^{-3}$ due to radial central cell transport and enhanced synchotron radiation. It is noteworthy that good performance is achievable here despite using generally more conservative parameters than the DT classical tandem proposed in Fowler et al. (2017) which had: $B_{m}=24~{}\mathrm{T}$, $a_{m}=0.21~{}\mathrm{m}$, $\langle n_{p}\rangle=2.6\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=55~{}\mathrm{m}$, $\beta_{c}\sim 1.0$, $n_{c}/n_{p}\sim 0.35$, $T_{i}=150~{}\mathrm{keV}$, and $T_{e}=150~{}\mathrm{keV}$.
@@
Figure 5: Contours of $T_{e}$ for operating points at given $\langle n\rangle$ and $E_{\mathrm{NBI}}$ satisfying the constraint equations (36). This plot uses the $a_{m}=0.15$ m POPCONs case found in Figure 4b. Black contour lines are fusion power in the central cell in MW.
@@
This POPCONs analysis exposed a number of surprising features of fusion relevant tandem mirror operation that were not previously well known. These discoveries are in part because, outside of the analysis in Fowler et al. (2017), axisymmetric tandem mirrors with high performance HTS magnets and high $R_{m}$ had never been considered. Older studies of tandem mirrors generally assumed end plug mirror ratios of order unity to account for minimum-B coil geometries and lower peak $B_{m}$ (Fowler & Logan, 1977). This analysis builds upon that in Fowler et al. (2017), and very importantly, it includes the effect of fusion alpha particle heating on the electron temperature. The key result here is that, when high-field mirror coils are available, classical tandem central cells can and should be operated at much higher density fractions $n_{c}/n_{p}$ than previously assumed. Much of the previous classical tandem mirror literature used $n_{c}/n_{p}=0.05-0.35$ (Dimov et al., 1976; Fowler & Logan, 1977; Hua & Fowler, 2004). The analysis here suggests this value should be higher, $n_{c}/n_{p}=0.35-0.6$. Strong alpha heating of the electrons obtained by, increasing the central cell density, boosting confinement with large $R_{m}$ from axisymmetrization, and directly heating the central cell ions with RF to enhance the fusion rate, allows the central cell to reach electron temperatures $\sim 2-3~{}T_{i}$ without supplemental electron heating as shown in Figure 5. This operational mode takes advantage of tandem mirror confinement (17) improving more rapidly with increasing temperature than it degrades with increasing the density ratio.
@@
In conclusion, this scoping shows that for high-field axisymmetric mirrors, classical tandem operation can reach appreciable Q and is a valid alternative to operation with thermal-barriers (Baldwin & Logan, 1979). These findings are in agreement with similar analyses by Pratt & Horton (2006) and Fowler et al. (2017). However, it has been found here that higher density operation than these previous analyses is possible thanks to alpha heating, somewhat relaxing the end plug performance criteria, but there are still some important caveats. The tandem mirror POPCON approach used here, much like the tokamak POPCON approach, is only able to identify operating points. It does not guarantee that an operating point it identifies is stable or achievable in actual operation. To determine this, a time dependent simulation which takes the mirror from an initial breakdown plasma to its final operating point must be performed. It is possible that, to actually reach the operational conditions proposed here, a transient heating source, such as a high-power electron cyclotron heating "sparkplug", might be needed to boost $T_{e}$ initially. This would drive up the end plug potentials and improve confinement enough for the central cell to generate sufficient alpha heating power to reach ignition. Furthermore, the effects of spatial gradients or turbulent transport were not included in this analysis. Generally, the spatial density gradients in tandem mirrors are predicted to be small (Mirin et al., 1983; Hua & Fowler, 2004), and the effects of electron temperature gradient turbulence, the main form of turbulent energy transport predicted to be present in tandem mirrors, are anticipated to be more tolerable than the temperature gradient driven turbulent transport in tokamaks (Hua & Fowler, 2004; Pratt & Horton, 2006). These topics will be treated in future work where gyrokinetic turbulence simulations at mirror central cell conditions will be performed to evaluate the growth rates of gradient driven instabilities. A time dependent solver based on those found in Mirin et al. (1983) and Hua & Fowler (2004) in order to evaluate the accessibility of these operating points under more realistic assumptions will also be implemented. In addition to these confinement issues, macrostability of the tandem system, specifically its stability against trapped particle modes and MHD, have not been addressed here and will be the topic of a future publication.
@@
For tandem mirror parameters similar to the ignited operating points in Figure 4b: $\tau_{c}=5~{}\mathrm{s}$, $\phi_{e}=400~{}\mathrm{keV}$, $\ell_{c}=50~{}\mathrm{m}$, $R_{mp}=8$, $R_{mc}=13.3$ (accounting for finite $\beta\sim 0.6$ in both mirror ratios), $\theta_{\mathrm{NBI}}=45^{\circ}$, $E_{i}=500~{}\mathrm{keV}$, $\mu=2.5~{}\mathrm{AMU}$, and $a_{m}\sqrt{R_{mp}}/\rho_{i0}=25$, it is calculated that $n_{c}/n_{p}>1000$ indicating DCLC likely cannot be stabilized with central cell outflow in the tandem mirrors under consideration here. As $\Gamma_{c,\mathrm{out}}$ is insufficient for DCLC stabilization, DCLC stability in Section 4 will be ensured by the suppression of the diamagnetic drift wave.
@@
4 Simulation of an end plug with optimized equilibrium, heating, and transport using RealTwin
The POPCON technique outlined in the previous section has defined what parameters the end plug must achieve in order to obtain a good central cell: the end plug mirror coils must have the highest field and bore possible, and an end plug must achieve the highest plasma density possible for the fewest watts of input power. In this section an end plug’s equilibrium transport will be optimized, first some assumptions will be used to simplify the problem and 0D physics and engineering constraints will be imposed on the end plug. The 0D constraints will be used to identify an end plug operating region. Then, simulations with the RealTwin will be performed to provide a coarse optimization of a simple mirror based on the 0D constraints to determine if the end plug conditions predicted by the 0D scoping can be obtained. Next, using the simple mirror density $\langle n_{p}\rangle$ parameters obtained by RealTwin, a new set of tandem POPCONs will be calculated. The range of tandem mirror ambipolar potentials and temperatures obtained from the POPCONs will be used to initialize a secondary set of RealTwin simulations with a fixed temperature electron backgrounds and an ambipolar potentials that correspond to the values obtained using POPCONs. These simulations will be iterated with the POPCONs until the end plug RealTwin simulation conditions and the POPCON conditions are self-consistent to provide a final equilibrium transport optimized end plug.
4.1 Assumptions and 0D constraints